syed98 (syedmohammed98):

If x^4 = 10 what is x^6? A) 10(square root 10) B) 100 C) 100(square root 10) D) 1000 E) 1000(square root 10)

OpenStudy (freckles):

hint: \[x^4=10 \\ x^\frac{4}{2}= 10^\frac{1}{2} \\ \]

OpenStudy (freckles):

so what is x^4 times x^2 =?

syed98 (syedmohammed98):

x^8?

OpenStudy (freckles):

it is x^6 since 4+2 is 6 not 8 but I was asking you for the other value like you can replace x^4 with 10 you can replace x^2 with?

OpenStudy (freckles):

use the hint above

OpenStudy (usukidoll):

it's like..... for example if we need to find x^8 and given x^4= 10 split the x^8 up with the exponent rule (x^4)(x^4) (10)(10) = 100... that's just an example

syed98 (syedmohammed98):

10(1/2)

OpenStudy (freckles):

10^(1/2) ? if so yes which is the same thing as saying sqrt(10)

syed98 (syedmohammed98):

so it's A?

OpenStudy (usukidoll):

so if \[x^2 = \sqrt{10} \] and \[x^4 = 10 \] what is \[x^6? \]

syed98 (syedmohammed98):

10(square root 10)

OpenStudy (freckles):

\[x^4=10 \\ x^4=10^{1} \\ \] \[x^\frac{4}{2}=10^\frac{1}{2} \\ \text{ so } x^2=10^\frac{1}{2} \\ \text{ and yes } \\ x^6=x^4 \cdot x^2 =10 \cdot \sqrt{10}\]

syed98 (syedmohammed98):

Thank you so much @UsukiDoll @freckles

OpenStudy (usukidoll):

it's like using the exponent rule to split up the x^6 \[x^6 \rightarrow x^{2+4} \rightarrow (x^2)(x^4)\] \[\sqrt{10}(10) \]

syed98 (syedmohammed98):

got it :) thnx

OpenStudy (usukidoll):

so there's a big difference between the addition/multiplication for the exponent rule and distributing the exponent 

syed98 (syedmohammed98):

i got it ...lol thnx :)

OpenStudy (usukidoll):